# A Cross-Cultural Analysis of Social Norms in Bollywood and Hollywood Movies

### v2

@@ 5.2 Cross-cultural variations in Social norms
Dataset of cross-cultural social norms
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (99.2%) shame-related dialogues out of 5363, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is expressed even without a reason to express affection, specifically toward close family members such as son. GPT-4 chat predicted a total of 10,445 social norms (See Table A2 for norm distribution). A sample of five reasons extracted for shame and pride is provided in Tables 5 and 6.
@@
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms.
5.2.1 Themes of Social Norms
We found twenty-five clusters for shame-related norms and twelve for pride-related norms using agglomerative clustering ( See Tables A12 and A13 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A10 and A11 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1.
@@ 5.2.1 Themes of Social Norms
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{Holly,Bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis).
@@
Poverty and inappropriate behaviors are more strongly associated with shame in Hollywood whereas gender conformity and disrespect are more strongly associated with shame in Bollywood (See Fig. 2). As expected, collectivist factors such as parent-related roles and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as incompetence and poor social etiquette evoke shame in Hollywood.
@@
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas serving the nation, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms.
5.2.2 Gender differences in Attribution of Self-conscious Emotions
For the Bollywood set, GPT-4 predicted 1632 targets as female and 4361 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A2), there are more male targets than females. This aligns with Geena Davis Inclusion Quotient (2016) findings that male characters get more screen time and speak more than female characters.
Using target gender as predicted by GPT-4 (See Table A2 for distribution), we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
@@ 5.2.2 Gender differences in Attribution of Self-conscious Emotions
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is similar (0.16 for Hollywood and 0.18 for Bollywood) irrespective of the movie industry.
Normative expectations evoking shame are surprisingly similar for females across movie industries (See Figs. 5 and 6). Sexuality (e.g., promiscuity, immodesty, gender roles) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. For pride, women express more pride for family roles. However, we do see a variation in themes for men-pride i.e., justice, bravery and achievement are major themes for pride in Bollywood whereas it is duty, and self-identity in Hollywood. We also note interesting reverse patterns e.g., self-identity is a source of pride for females in Bollywood whereas it is for males in Hollywood.
@@ 6 Discussion
Cross-cultural variations in expressions of moral emotions
@@
The entailment of shame tells us about undesired behavior, whereas pride reveals desired behaviors in a society. While self-conscious emotions are culture-agnostic markers of unspoken social rules, the expressions of these emotions vary across cultures. We uncovered these variations in Sec 5.1. Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe a strong correlation of first-person singular pronouns with shame and its infrequent use in Hollywood movies compared to Bollywood movies (See Table 2). In contrast, shame is interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) nature of shame in Bollywood movies (See Table A6). Moreover, the contrasting tenses coupled with different emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). The correlation of shame with categories such as sexual, female, conflict, swear is exclusive to Bollywood, likely due to emphasis on conformity in Indian society. Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial).
Moral Emotion-based Prompting
Compared to GoEmotions dataset Demszky et al. (2020) having 817 samples for embarrassment and 452 samples for pride, we provide a much larger cross-cultural dataset of shame (6583 samples) and pride (5804 samples) related discourse, collected from movie subtitles across two cultures. Additionally, we release 10K culture-specific social norms. These can be used for downstream applications like evaluation of culturally aware language inference and achieving pluralistic values in LLMs.
@@
Self-conscious emotion-based prompting helped in mitigating cultural bias in LLMs-generated text and potential stereotyping during norm discovery; as evident from the specificity of the norms extracted using our approach ( See Tables 5 and 6). More importantly, these norms reflect the cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter vs. honesty in business, returning calls after date night as in Table 5) revealing the differences in socially expected behaviors in Indian society and the U.S. society. The excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
Cross-cultural variations in Social Norms
@@
To tease apart underlying values evoking moral emotions across cultures, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (i.e., nation honor, family honor) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes still reflect the patterns seen during LIWC analysis, supporting the use of moral emotions for overcoming cultural bias in LLMs.
Gender Differences
@@
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. Surprisingly, the gender gap in pride- and shame-related dialogues is similar across both societies. While our initial hypothesis that "more shame is attributed to women in Indian society compared to American society" (0.18 in Bollywood vs 0.16 in Hollywood) is true, the difference is negligible. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
@@ 7 Conclusion
The prompts ("You study Indian society" vs "You study Western society" described in Tables A4 and A5 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
@@ A.1 Search Keywords
A.2 Norms Distributions
@@
male 3102 1259 591 776
female 1306 326 246 236
not known 913 652 319 719
@@ A.3 Annotation
A.4 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A4 and A5.
@@ A.4 Prompts
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud. If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
@@
Output: mother, male, provide care for old parents
Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre
Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming
The answer should be short and in CSV format. $<$ shamed, gender, reason$>$
Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin.
Output: girl’s parent, NA, not able to marry off their daughters
Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’
Output: System, NA, poor law and regulations"
Table A4: Bollywood: Prompts for norm discovery using GPT-4 Chat.
@@ 2. Identify the primary reason for shaming
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender.
2. Identify the reason behind feeling proud? If not explicit, return NA.
The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir?
Output: Sir, male, winning olympic gold
Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart.
Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify
1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender.
2. Identify the primary reason for shaming.
The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$
Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake.
Output: NA, male, not returning borrowed money
Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours.
Output: Snake, male, stealing and preying on people"
@@
Table A5: Hollywood: Prompts for norm discovery using GPT-4 Chat.
A.5 LIWC Correlation Results

### v3

@@ 4.1 How do expressions of shame/pride vary across cultures?
4.2 How do reasons behind shame/pride vary across cultures?
Cross-Cultural Norms Dataset
For the Bollywood set, GPT-4 chat predicted reasons (also implicit social norms) for 5321 (98.4%) shame-related dialogues out of 5409, and 2237 (74.6%) pride-related dialogues out of 2999. For the Hollywood set, GPT-4 chat predicted a reason for 1156 (94.6%) shame-related dialogues out of 1221 and 1731 (61.7%) pride-related dialogues out of 2805. Upon manual analysis, we found that pride is also used to express affection, specifically toward close family members such as sons - therefore, a lower number of norms is associated with pride compared to shame. GPT-4 chat predicted a total of 10,445 social norms (See Table A3 for norm distribution). Prompting "reasons" behind shame and pride allowed us to capture high specificity in cultural norms (See Tables 4 and 6 for examples).
4.2.1 Manual Evaluation
Two volunteers manually verified the predicted gender for "the person experiencing shame/pride" and the predicted "reason" in a randomly sampled set of $100$ dialogues each from Hollywood and Bollywood. The annotator who labeled the Bollywood set is an Indian, aware of social roles and expectations in Indian society. Likewise, the annotator for the Hollywood set is an American with a nuanced understanding of the social norms of the U.S.A.
For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table 5). There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
Gender Evaluation
Incorrect Ambiguous
Bollywood 8 15 (3 Female)
Hollywood 5 10 (3 Female)
Social Norms/Reason Evaluation
Incorrect Ambiguous
Bollywood 11 9
Hollywood 2 1
Table 5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
@@ 4.2.1 Manual Evaluation
4.2.2 Thematic Analysis
Twenty-five clusters for shame-related norms and 13 for pride-related norms were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
@@ 4.2.2 Thematic Analysis
Here, $\mathcal{D}_{industry_{t_{i}}}$ represents dialogues to $industry\in\{holly,bolly\}$ and ${t_{i}}\in themes$ (as depicted on Y-axis in Fig 2 and 3).
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates insignificant difference i.e., $p>0.05$
@@
Themes such as lack of accountability and poverty are more common in Hollywood whereas norms related to sexual behavior and gender roles are more prevalent in Bollywood (See Fig. 2). As expected, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, poor social etiquette and incompetence evoke shame in Hollywood.
@@
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas Ethnolinguistic identity, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates insignificant difference i.e., $p>0.05$
4.2.3 Gender differences in Social Emotions
For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A3), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
Using predicted gender, we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
@@ 4.2.3 Gender differences in Social Emotions
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Normative expectations evoking shame are similar for females across movie industries (See Fig. A1). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A2).
@@ 5 Discussion
Detecting social emotions
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related discourse with over 10K social norms. To the best of our knowledge, the only other corpus having samples related to social emotions is GoEmotions dataset Demszky et al. (2020) with 817 samples for embarrassment and 452 samples for pride. Detecting social emotions, and recognizing how appraisals and expressions of such emotions vary will improve human-computer communications, specially in domains of AI powered mental health therapy Kim et al. (2011), cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
Cross-cultural variations in expressions of social emotions
@@
Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe its infrequent use in Hollywood movies and self-focus (reflecting internal shame) (See Table A2 and 2). In contrast, shame is considered interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) or public nature of shame in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10).
Cross-cultural variations in social sanctions and approvals
@@
Eliciting reasons leading to social emotions shame and pride revealed socially sanctioned behaviors across cultures. We observed high specificity in extracted social norms revealing cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter in Bollywood vs. honesty in business, returning calls after date night in Hollywood as in Table 4). We acknowledge that the excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
@@
To tease apart ways in which cultures differ, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (e.g., Ethnolinguistic identity) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are very similar to patterns seen during LIWC analysis, supporting the use of social emotions for overcoming cultural bias in LLMs.
Women face more social sanctions than men
@@
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
6 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023).
@@ 7 Conclusion
The prompts ("You study Indian society" vs "You study Western society" described in Tables A5 and A6 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in writing and refining this manuscript.
@@ A.1 Search Keywords
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table A2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
A.2 Examining Implicit shame:
A.3 Norms Distributions
@@
male 2690 1259 591 776
female 1215 326 246 236
Not clear 1416 652 319 719
@@ A.4 Annotation
A.5 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A5 and A6.
@@ A.5 Prompts
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud. If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$
@@
Output: mother, male, provide care for old parents Input: I’ve heard there’s a promising young student in your school What’s his name? He’s made us proud in long jump, we are here to felicitate him Call him Show yourself, Raju Tempre Output: authority, NA, Sports achievement"
"You are an experienced social scientist and you study Indian society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming The answer should be short and in CSV format. $<$ shamed, gender, reason$>$ Input: And should we bow before others begging....them to marry our daughters? This shall not happen. Neither will the girls be alive here nor shall....we be ashamed of ourselves. You cannot kill the life which God has given. I won’t let you commit the sin. Output: girl’s parent, NA, not able to marry off their daughters Input: Black marketers are now in the open. And the thieves too Politics is in a great mess Shame on this system. There’s no democracy Get rid of these politicians The gong has struck..’’Our hearts are swaying to it’s beats’’ Output: System, NA, poor law and regulations"
Table A5: Bollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
@@
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud? If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir? Output: Sir, male, winning olympic gold Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart. Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
@@
Table A6: Hollywood: Prompts for norm discovery using GPT-4 Chat. The temperature was set to 00 to minimize randomness.
A.6 LIWC Correlation Results